OpenStudy (anonymous):

give me conceptt nd defination of domain and range in two variable function ?????????

OpenStudy (anonymous):

Domain means the set of all possible x values that would make sense. ie. equation y=x+2. x can take on any value. y=square root x, x can only be 0 or all positive numbers. Range means the set of all y values that could be generated if you used all those possible x values. ie. y=x+2, the range is all real numbers. y = square root x, the range is 0 and all positive numbers.

OpenStudy (anonymous):

what is domain of f(x,y)=ln(x2-y2) ?

OpenStudy (anonymous):

???????????? helo

OpenStudy (dan815):

what does a ln graph look like

OpenStudy (anonymous):

no idea ?

OpenStudy (dan815):

go find out that first, and tell me where ln graph is undefined

OpenStudy (anonymous):

okx

OpenStudy (anonymous):

defined at ln(x)>0

OpenStudy (dan815):

ok good

OpenStudy (dan815):

so ln(something) where something has to be more than 0

OpenStudy (dan815):

now your something is x^2-y^2

OpenStudy (anonymous):

ln(1)=0 ln(0)=\[ \infty\]

OpenStudy (anonymous):

means greater than 0?

OpenStudy (dan815):